I'm replacing the tires on my truck and want to know my engine speed (RPM) with the new tires. Old tire revolutions per mile = 794 / new tire revolutions per mile = 782. Engine speed at 65MPH with old tires = 3,100 RPM. What is my engine speed going to be with the new tires?
Thanks,
Mike

Answer by checkley71(8403)

792/782=X/65 NOW CROSS MULTIPLY
782X=792*65
782X=51480
X=51480/782
X=65.83 THE NEW SPEED BECAUSE OF THE LARGER TIRES.